OpenStudy (anonymous):

The west side factory produces 450 boxes of cereal every day. A machine drops a single toy car into each box of cereal. Today the machine is loaded with 115 blue cars, 190 green cars and 145 red cars. what is the probability that one of the boxes of cereal has a blue car inside? explain.

OpenStudy (anonymous):

@Directrix

OpenStudy (campbell_st):

P(blue) = total number of blue ------------------ total number of boxes

OpenStudy (anonymous):

so do i divide each number by the total number of boxes?

OpenStudy (anonymous):

115/450 then 190/450, 145/450

OpenStudy (campbell_st):

no, you just want the porbability of a blue toy.. \[P(blue) = \frac{number~ of ~blue~toys}{total ~number ~ of ~boxes}\]

OpenStudy (anonymous):

oh okay hold on

OpenStudy (anonymous):

0.255555556

OpenStudy (anonymous):

@mathmath333

OpenStudy (campbell_st):

well I would write the answer as a simplified fraction so P(blue) = 28/90 but otherwise you're corrrect

OpenStudy (anonymous):

so how do i get it into a fraction?

OpenStudy (anonymous):

it said that was wrong......

OpenStudy (mathmath333):

try this \(\dfrac{28}{90}=\dfrac{14}{45}\)

OpenStudy (anonymous):

well ive already submitted.....

OpenStudy (campbell_st):

well I hope it has the decimal in the answer list

OpenStudy (mathmath333):

ok